# A Survey of RWKV

### v1

@@ 1 Introduction
Others[105] [106] [107] [108] [109] [110] [111]] ] [Natural Language
@@
RWKV-TS [130] / TIM [131] / RRWKV [132] / LKPNR [133] / Music-Diff [134]
@@
Diffusion-RWKV [158] / RWKV-IR [159]
Others [160] [161]] ] [Web
@@
Others [197] [198] [199] [200] [201] [202] [203] [204] ] ] [Others,text width=3em [Voice Activity Detection [205] / RWKV-Transducer [206] / RWKV-ASR [207]
Spiking Mixer [208] / MiniRWKV-4 [209] / RWKV-RAG [100]
@@
AudioRWKV [213] / RWKV-Reward-Enhanced [214] / RWKV-MusicGenerator [215]
Music-Genre-Classification-RWKV [216] / RWKV-Twitter-Bot-Detection [217]
@@
ChatRWKV-PC [256] / Generative Calibration [257] / PlantFlowerDatasets [258]] ] ] [Future Directions and
@@ 3.5 Various Implementations
The RWKV authors offer an official PyTorch implementation as detailed in version [44], encompassing the implementation specifics for RWKV-1 through RWKV-6. Their efforts are further extended with the ongoing development of a laboratory version of RWKV-7. Additionally, they introduced a Chatbot framework called ChatRWKV [45], designed to enable users to construct their own RWKV inference engines.
@@ 4 Applications of the RWKV Model
The RWKV model integrates the advantages of recurrent mechanisms with those of attention-based approaches. This unique architectural design enhances its efficiency and scalability, enabling it to adeptly tackle a wide range of problems. This chapter provides an overview of the practical applications of the RWKV model, focusing on its deployment in areas like natural language generation, natural language understanding, computer vision, and web-related tasks. Figure 3 illustrates examples of downstream tasks that make use of RWKV-based models. Additionally, the discussion extends to the use of RWKV in other NLP tasks, where its performance is evaluated across diverse assessment frameworks.
@@ 4.1 Natural Language Generation
Within interactive settings, RWKV is used for role-playing scenarios [103, 104], and also for generating WeChat responses via projects like Chatrwkv [101, 102]. Furthermore, RWKV is implemented in a Telegram LLM bot written in Rust, which can operate using either RWKV or Llama2 models [105]. Chatbots with RWKV capability also facilitate multimedia interactions, such as sending voice messages and stickers [106, 107]. Additionally, chatbots using the Rasa NLU framework alongside RWKV-4-Raven can address frequently asked questions in French, English, and Tunisian Arabic [108]. RWKV extends its versatility to generating MIDI music [109, 110]. Finally, its creative versatility is showcased in a mini-game that fuses Stable Diffusion with the RWKV music model [111]. Across these diverse applications, RWKV models manifest substantial potential in advancing natural language generation and refining human-computer interaction.
@@ 4.3 Other NLP tasks
The Retrospected Receptance Weighted Key Value (RRWKV) framework incorporates retrospective features that improve its capacity to understand long-term dependencies, thereby overcoming the challenges faced by RWKV in processing distant information [132]. In addition, the application of transfer learning techniques significantly boosts the efficiency of domain generation algorithm classifiers, which facilitates accurate distinction between non-malicious and harmful exemplars [141]. In the realm of personalized news recommendation systems, the fusion of large language models with knowledge graphs significantly boosts the precision of recommendations [133]. The Decision-RWKV model, which blends decision transformers with experience replay frameworks, exhibits strong performance in reinforcement learning and decision-making tasks [135]. To combat the forgetting phenomenon prevalent in language models, a strategy utilizing synthetic gradients is introduced. This method encodes prompts directly, thereby enhancing memory retention during text generation [136]. Leveraging the RWKV architecture, the SpikeGPT model employs event-driven sparse activations to generate natural language effectively [137]. Within the domain of music generation, the Music-Diff architecture improves the diversity of samples by employing a joint probability diffusion model. This approach effectively overcomes the shortcomings of conventional language models typically used in generative tasks [134].
@@ 4.4 Computer Vision
The LSTM CrossRWKV (LCR) framework employs a Cross RWKV gate along with an improved LSTM mechanism to adeptly capture both spatial and temporal aspects pertinent to video analysis [154]. For image-text interactions, RWKV-CLIP stands out as the pioneer RWKV-integrated visual-language representation model, merging the parallel training efficiency of transformers with the inference efficacy of RNNs [155]. Studies involving Spiking Neural Networks (SNNs) underscore their capabilities in generating images [156]. Moreover, approaches for predicting multimodal pedestrian trajectories utilize conditional variational autoencoders to adeptly account for the diversity and uncertainty in human motion [157].
@@ 4.7 Others
One major breakthrough in technology is the advancement of speech recognition systems. Notably, RWKV-based semantic voice activity detection (VAD) has shown enhanced resilience to noise interference [205]. Additionally, RWKV has been effectively utilized in the realm of streaming Automatic Speech Recognition (ASR). Models such as the RWKV-Transducer have achieved competitive levels of accuracy while simultaneously optimizing latency and reducing memory consumption [206, 207].
@@
A method for image-text alignment employing a frozen RWKV-4-World-0.4B model highlights its potential in multimodal tasks [211]. The breadth of RWKV’s capabilities encompasses a wide array of tasks, including sequence classification, embedding generation, and cross-encoder applications, demonstrating its versatility in tackling various challenges in machine learning [212]. Moreover, RWKV has been successfully applied in audio pattern recognition [213], as well as in the training of reward models enhanced by human feedback within the context of reinforcement learning [214]. Its applications further extend to music generation [215], music genre classification [216], and the identification of Twitter users who are bots [217].
@@
The notable performance of RWKV across various applications is worth highlighting. Notable implementations involve the utilization of the GGML library for offline operations on iOS and macOS, as detailed in [247], as well as prompt projects that integrate RWKV with Langchain, referenced in [248]. Additionally, the visualization service created for the RWKV-Ouroboros project, cited in [249], exemplifies the vast applicability of RWKV. Moreover, deployment scripts for the automatic deployment of GGML/GGUF models on Android have been developed, as seen in [250]. The integration of RWKV with the Llama architecture [251] and the implementation of GPTQ quantization techniques [252] further underscore the innovative progress within this area.
Transferring RWKV-LM models to the ROCm platform, as outlined in [253], employing inference through OpenGL technology [254], and merging RWKV with server-based applications designed for question-answering tasks [255] create promising new opportunities for discovery. The PC version of ChatRWKV [256] is equipped with a user-friendly interface, enhancing accessibility. Moreover, the exploration into generative calibration methods for RWKV, as discussed in [257], focuses on strategies for enhancing model performance during context learning. The availability of open datasets for RWKV [258] provides crucial resources supporting researchers in their work.

### v2

@@ 1 Introduction
Others[105] [106] [107] [108]] ] [Natural Language
@@
RWKV-TS [127] / TIM [128] / RRWKV [129] / LKPNR [130]
@@
Diffusion-RWKV [154] / RWKV-IR [155] / StyleRWKV [156]
Others [157] [158]] ] [Audio and
Music Tasks,text width=5em [Android RWKV MIDI [159] / RWKV-v4 MIDI [160] / RWKV-v4 music models  [161]
Music-Diff [162] / Voice Activity Detection [163] / RWKV-Transducer [164]
RWKV-ASR [165] / AudioRWKV [166] / RWKV-MusicGenerator [167]
Music-Genre-Classification-RWKV [168]  RWKV-SpeechChat [169]] ] [Web
@@
Others [205] [206] [207] [208] [209] [210] [211] [212] ] ] [Others,text width=3em [Spiking Mixer [213] / MiniRWKV-4 [214] / RWKV-RAG [100]
@@
RWKV-Reward-Enhanced [218] / RWKV-Twitter-Bot-Detection [219]
@@
ChatRWKV-PC [258] / Generative Calibration [259] / RWKV- DPA [260]
RWKV-edge [261] / PCF-RWKV [262] / L3TC [263] /PlantFlowerDatasets [264]] ] ] [Future Directions and
@@ 3 RWKV
Currently, the official versions of RWKV mainly include three types: RWKV-4, RWKV-5 (Eagle), and RWKV-6 (Finch). Among them, the commonly used version is RWKV-4. In addition, the latest RWKV-7 (Goose) has released a preview version and the corresponding code, but the specific document has not yet been made public yet [^1] https://www.rwkv.com .
@@ 3.5 Various Implementations
The RWKV authors offer an official PyTorch implementation as detailed in version [44], encompassing the implementation specifics for RWKV-1 through RWKV-6. Their efforts are further extended with the ongoing development of a laboratory version of RWKV-7 (Goose). Additionally, they introduced a Chatbot framework called ChatRWKV [45], designed to enable users to construct their own RWKV inference engines.
@@ 4 Applications of the RWKV Model
The RWKV model integrates the advantages of recurrent mechanisms with those of attention-based approaches. This unique architectural design enhances its efficiency and scalability, enabling it to adeptly tackle a wide range of problems. This chapter provides an overview of the practical applications of the RWKV model, focusing on its deployment in areas like natural language generation, natural language understanding, computer vision, audio and music tasks, and web-related tasks. Figure 3 illustrates examples of downstream tasks that make use of RWKV-based models. Additionally, the discussion extends to the use of RWKV in other NLP tasks, where its performance is evaluated across diverse assessment frameworks.
@@ 4.1 Natural Language Generation
Within interactive settings, RWKV is used for role-playing scenarios [103, 104], and also for generating WeChat responses via projects like Chatrwkv [101, 102]. Furthermore, RWKV is implemented in a Telegram LLM bot written in Rust, which can operate using either RWKV or Llama2 models [105]. Chatbots with RWKV capability also facilitate multimedia interactions, such as sending voice messages and stickers [106, 107]. Additionally, chatbots using the Rasa NLU framework alongside RWKV-4-Raven can address frequently asked questions in French, English, and Tunisian Arabic [108]. Across these diverse applications, RWKV models manifest substantial potential in advancing natural language generation and refining human-computer interaction.
@@ 4.3 Other NLP tasks
The Retrospected Receptance Weighted Key Value (RRWKV) framework incorporates retrospective features that improve its capacity to understand long-term dependencies, thereby overcoming the challenges faced by RWKV in processing distant information [129]. In addition, the application of transfer learning techniques significantly boosts the efficiency of domain generation algorithm classifiers, which facilitates accurate distinction between non-malicious and harmful exemplars [137]. In the realm of personalized news recommendation systems, the fusion of large language models with knowledge graphs significantly boosts the precision of recommendations [130]. The Decision-RWKV model, which blends decision transformers with experience replay frameworks, exhibits strong performance in reinforcement learning and decision-making tasks [131]. To combat the forgetting phenomenon prevalent in language models, a strategy utilizing synthetic gradients is introduced. This method encodes prompts directly, thereby enhancing memory retention during text generation [132]. Leveraging the RWKV architecture, the SpikeGPT model employs event-driven sparse activations to generate natural language effectively [133].
@@ 4.4 Computer Vision
The LSTM CrossRWKV (LCR) framework employs a Cross RWKV gate along with an improved LSTM mechanism to adeptly capture both spatial and temporal aspects pertinent to video analysis [150]. For image-text interactions, RWKV-CLIP stands out as the pioneer RWKV-integrated visual-language representation model, merging the parallel training efficiency of transformers with the inference efficacy of RNNs [151]. Studies involving Spiking Neural Networks (SNNs) underscore their capabilities in generating images [152]. Moreover, approaches for predicting multimodal pedestrian trajectories utilize conditional variational autoencoders to adeptly account for the diversity and uncertainty in human motion [153]. StyleRWKV [156] leverages RWKV models for efficient style transfer, introducing Re-WKV attention, a Deform-Shifting layer, and Skip Scanning to enhance global and local context while maintaining low memory usage and linear time complexity.
4.5 Audio and Music Tasks
In music generation, the Music-Diff architecture enhances sample diversity by using a joint probability diffusion model, overcoming the limitations of traditional language models in generative tasks [162]. RWKV has extended its capabilities to MIDI music generation [159, 160] and has been applied in music genre classification [168]. Its creative versatility is further demonstrated in a mini-game that integrates Stable Diffusion with the RWKV music model [161].
In the field of speech recognition, RWKV has made notable advancements. RWKV-based semantic voice activity detection (VAD) shows enhanced resilience to noise interference [163]. At the same time, the RWKV-Transducer model has achieved competitive accuracy in streaming Automatic Speech Recognition (ASR), optimizing both latency and memory consumption [164, 165]. Additionally, RWKV-SpeechChat uses the frozen 3B RWKV model to support real-time conversational tasks, including ASR, speech translation, and speech question answering [169]. RWKV has also been successfully applied in audio pattern recognition [166].
@@ 4.8 Others
A method for image-text alignment employing a frozen RWKV-4-World-0.4B model highlights its potential in multimodal tasks [216]. The breadth of RWKV’s capabilities encompasses a wide array of tasks, including sequence classification, embedding generation, and cross-encoder applications, demonstrating its versatility in tackling various challenges in machine learning [217]. Moreover, RWKV has been successfully applied in the training of reward models enhanced by human feedback within the context of reinforcement learning [218]. Its applications further extend to the identification of Twitter users who are bots [219].
@@
The notable performance of RWKV across various applications is worth highlighting. Notable implementations involve the utilization of the GGML library for offline operations on iOS and macOS, as detailed in [249], as well as prompt projects that integrate RWKV with Langchain, referenced in [250]. Additionally, the visualization service created for the RWKV-Ouroboros project, cited in [251], exemplifies the vast applicability of RWKV. Moreover, deployment scripts for the automatic deployment of GGML/GGUF models on Android have been developed, as seen in [252]. Compression techniques specifically designed for RWKV, such as low-rank approximation, sparsity predictors, and clustering heads, have achieved significant results, enabling up to 4.95–3.8x compression with minimal accuracy loss [261]. The integration of RWKV with the Llama architecture [253] and the implementation of GPTQ quantization techniques [254] further underscore the innovative progress within this area.
Transferring RWKV-LM models to the ROCm platform, as outlined in [255], employing inference through OpenGL technology [256], and merging RWKV with server-based applications designed for question-answering tasks [257] create promising new opportunities for discovery. The PC version of ChatRWKV [258] is equipped with a user-friendly interface, enhancing accessibility. Moreover, the exploration into generative calibration methods for RWKV, as discussed in [259], focuses on strategies for enhancing model performance during context learning. In V2X environments, an RWKV-DPA [260] estimator was proposed, combining RWKV network and data-pilot aided estimation. It outperformed other DL-based estimators under IEEE 802.11p standard with lower complexity. The availability of open datasets for RWKV [264] provides crucial resources supporting researchers in their work. The PCF-RWKV [262] model leverages RWKV architecture, task-specialized LoRAs, and Multi-Agents technology to automate lifecycle inventory construction and carbon footprint calculation. The L3TC [263] method integrates RWKV models, an outlier-aware tokenizer, and high-rank reparameterization to enable efficient, low-complexity text compression.